# TimeViper: A Hybrid Mamba-Transformer Vision-Language Model for Efficient Long Video Understanding

### v1

@@ Abstract
00footnotetext: $\dagger$ Corresponding author: Qin Jin; Equal contribution; $\ddagger$ Project lead
@@ 3.2 Token Information Transfer
TimeViper (ours w/ TransV) Nanov2-9B [basant2025nanov2] 56.2 52.0 63.1 56.9 48.2 35.6 37.9 38.7
3.3 Training Procedure
4 Experiments
We first describe the training and evaluation setups in Section 4.1, followed by the main results and ablation studies in Section 4.2. Finally, in Section 4.3, we provide a qualitative analysis illustrating how hybrid models and Transformers interpret visual content through attention visualizations.
4.1 Experimental Setup
@@
Table 2: Ablation of TransV. The “uni_7_0.5-attn_39_0.9” denotes applying uniform TranV at both the 7th layer with a dropping rate of $p=50\%$ and attention-guided TransV at the 39th layer with $p=90\%$. “TDuni” denotes dropping tokens uniformly.
LLM Compression max frame VideoMME VDC Charades
1 none 5k 58.8 39.7 40.5
2 TDuni_7_0.5 8k 57.3 39.0 26.1
3 uni_7_0.5 8k 56.7 38.9 38.1
4 uni_7_0.5-uni_39_0.9 >>10k 56.2 39.1 37.9
5 uni_7_0.5-attn_39_0.9 >>10k 56.6 39.0 37.9
@@ 4.2 Main Results
Ablation choices of TransV. We investigate whether and how TransV should be applied in shallow and deep layers, and how much it can expand the number of supported frames. As shown in Table 2, the baseline can process about 5K input frames. By comparing rows 1, 2 and 3, after introducing token dropping or token compression, the model becomes capable of handling over 8K frames. Comparing rows 2 and 3, using TransV effectively compensates for the Charades performance drop from 26.1 to 38.1, indicating that it successfully achieves token transfer. Comparing rows 4 and 5, among the approaches for further compressing information in deeper layers, attention-guided TransV yields higher MCQ performance, while the differences on VDC and Charades are marginal. Furthermore, transferring token information in deeper layers allows the model to handle tens of thousands of frames with only a 0.1 point drop on VideoMME comparing rows 3 and 5.
4.3 Qualitative Analysis
To better understand how hybrid MLLMs differ from Transformer-based MLLMs in processing multimodal inputs, we analyze attention behaviors across layers. For Mamba layers, we follow [ameen2025mambavis] to define the attention pattern, while for Transformer layers, we use the attention weights.
@@ 5 Conclusion
This work takes an initial step toward understanding and compressing hybrid vision-language models for long videos. Our analysis shows that, in hybrid MLLMs, visual information gradually shifts into text tokens as depth of LLM layer increases, resulting in substantial redundancy among vision tokens in deep layers. Motivated by this observation, we introduce TimeViper, a Mamba-Transformer hybrid model equipped with TransV, an internal LLM token transfer module that compresses vision information into text tokens. This design enables efficient processing of hour-long videos while maintaining strong multimodal understanding. TimeViper achieves promising performance for long video understanding across multi-choice video QA, temporal video grounding, and detailed video captioning.
@@
Limitations. First, our current performance still falls short of the SOTA models due to limited training data and insufficient model training. Second, while TransV enables processing over 10,000 frames, the model has not been trained on videos of such duration.

### v2

@@ Abstract
00footnotetext: $\dagger$ Corresponding author; $\ddagger$ Project lead; Equal contribution.
spacThis work was done during their internship at Xiaomi.
@@ 3.2 Token Information Transfer
3.3 Training Procedure
4 Experiments
We first describe the experimental setups in Section 4.1, followed by the ablation studies in Section 4.2 and main results in Section 4.3. Finally, in Section 4.4, we provide a qualitative analysis illustrating how hybrid models and Transformers interpret visual content through attention visualizations.
4.1 Experimental Setup
Table 1: Ablation of TransV choices. The “uni_7_0.5-attn_39_0.9” denotes applying uniform TranV at the 7th layer with a dropping rate of $p=50\%$ and attention-guided TransV at the 39th layer with $p=90\%$. “TDuni” denotes uniform token dropping.
Method max frame VideoMME VDC Charades
1 none 5k 58.8 39.8 40.5
2 TDuni_7_0.5 8k 57.3 39.0 26.1
3 uni_7_0.5 8k 56.7 38.9 38.1
4 uni_2_0.5 9k 56.1 39.7 38.2
5 uni_7_0.9 >>10k 53.4 37.9 34.6
6 uni_7_0.5-uni_39_0.9 >>10k 56.2 39.4 37.9
7 uni_7_0.5-attn_39_0.9 >>10k 56.6 39.1 37.9
Figure 5: Comparison of GPU memory usage during inference. While ToMe extends the context window to about 5K frames, TransV efficiently scales beyond 10K frames.
Figure 6: Comparison of prefilling time. TransV incurs no additional latency at low frame inputs (e.g., 64 frames) while significantly reducing prefilling time at high frame inputs. For instance, at 4,096 frames, TransV reduces prefilling time by 15.7% compared to the ToMe baseline.
4.2 Ablation Study
For representation simplicity, each TransV is denoted as “type_layer_ratio-…”, for example, “uni_7_0.5-attn_39_0.9” represents applying uniform TransV in the 7th layer with a ratio of 50% and attention-guided TransV in the 39th layer with ratio of 90%.
Impact of compression components on GPU memory consumption. We apply ToMe in the projector and TransV in the LLM. Benefiting from the hybrid Mamba-Transformer backbone, both memory usage and prefilling time generally grow approximately linearly with input length. Figure 5 reports memory usage as the number of frames increases. The vanilla model runs “Out of Memory” error at merely 128 frames. TimeViper applying ToMe in the projector alleviates the initial token burden, extending the limit to approximately 5K frames. TimeViper with TransV further enables better scalability: at 4,096 frames, it reduces memory consumption by 54.8% compared to TimeViper, and can handle 10K+ frames with ample margin. This highlights the complementary roles of token compression in the projector and within the LLM.
Impact of compression components on prefilling time. As shown in Figure 6,the vanilla model already incurs 4.5s latency at 64 frames. TimeViper drastically reduces this to 0.4s, and TransV further decreases prefilling time, with the effect becoming more pronounced as the number of frames increases. Notably, at 4,096 frames, TransV reduces prefilling time by 15.7% compared to TimeViper.
@@ 4.2 Ablation Study
TimeViper (ours w/ TransV) Nanov2-9B [basant2025nanov2] 56.2 52.0 63.1 56.9 48.2 35.6 37.9 39.1
Figure 7: Comparison of performance as the number of input frames increases on long-video understanding benchmarks. We train our models with 256 frames as inputs, and sample 1 frame per second during evaluation. The x-axis here denotes the maximum number of frames. If a video exceeds this length, we take only the first max frames for inference.
TransV placement in shallow layers. As shown in Table 1, the TimeViper baseline can process approximately 5K input frames. By comparing rows 1, 2, and 3, we observe that introducing token dropping or token compression enables the model to handle over 8K frames. Comparing rows 2 and 3, using TransV effectively mitigates the Charades performance drop, from 26.1 to 38.1, indicating that it successfully facilitates token transfer. Comparing rows 3 and 4, compressing at the 7th layer does not necessarily outperform compression at the 2nd layer on the VDC or TVG benchmarks. For example, compression at the 7th layer outperforms the 2nd layer by 0.6 points on MCQ, but performs 0.8 points worse on VDC.
TransV placement in deep layers. From rows 6 and 7 in Table 1, attention-guided TransV yields higher MCQ performance of 56.6 than uniform TransV’s 56.2, with minor differences on VDC and Charades. Moreover, transferring token information in deeper layers significantly increases the model’s long-context capacity: comparing rows 1 and 3, the model handles tens of thousands of frames with only a 0.1 drop on VideoMME.
Compression rate for TransV in shallow layers. We evaluate a higher compression rate of $p=90\%$ at the 7th layer in Table 1. Larger compression rate allows the model to process more frames, but it comes with a significant performance drop. Comparing rows 4 and 5, after increasing compression rate from $50\%$ to $90\%$, the accuracy on VideoMME decreases from 56.7 to 53.4.
@@ 4.3 Main Results
Effect of increasing the number of inference frames. Since the model is trained with 256 frames, we evaluate test-time scalability by varying the number of input frames. As shown in Figure 7, TimeViper scales robustly with longer contexts across four long video understanding benchmarks. For example, when increasing the input frames from 256 to 512 frames, MLVU improves from 65.64 to 69.00, and LVBench increases from 35.53 to 37.0.
4.4 Qualitative Analysis
To better understand how hybrid MLLMs differ from Transformer-based MLLMs in processing multimodal inputs, we first formalize the definitions of attention scores used in both Mamba-2 and self-attention layers and then analyze attention behaviors across layers. For Mamba layers, we follow [ameen2025mambavis] to define the attention pattern, while for Transformer layers, we use the attention weights.
@@ 4.4 Qualitative Analysis
Definition of attention score. For self-attention (Equation 2), the attention score $M_{j,i}\in\mathbb{R}$ from $x_{j}$ to $x_{i}$ is:
$y_{i}=\mathrm{Softmax}(\frac{Q_{i}K_{\leq i}^{T}}{\sqrt{D}})\cdot V_{\leq i}=\sum_{j=1}^{i}M_{i,j}V_{j}$ (7)
For the SSM block, we rewrite Equation 1 to express its attention pattern as the weighted sum from inputs $x_{1},\dots,x_{i}]$ to the output $y_{i}$:
$y_{i}=\sum_{j=1}^{i}C_{i}^{T}(\prod_{k=j+1}^{i}A_{k})B_{j}x_{j}=\sum_{j=1}^{i}M^{\prime}_{i,j}x_{j}$ (8)
Here, $|M^{\prime}_{i,j}|\in\mathbb{R}^{+}$ serves as the “attention score” [ameen2025mambavis, zimerman2025mambavis] from $x_{j}$ to $x_{i}$ within the SSM block. Although both the self-attention and Mamba mechanisms employ a multi-head design [vaswani2017attention, dao2024mamba2] along the hidden dimension, we omit this detail in the equations for simplicity.
@@
Decrease of attention to vision tokens across layers. Comparing the second row with the first row in Figure 8, we observe a clear downward trend in attention scores assigned to vision tokens as the layers deepen. This phenomenon is consistent with our broader findings: as the model processes more layers, visual information becomes increasingly redundant and is subsequently deprioritized by instruction and response tokens.
@@ 5 Conclusion
This work takes an initial step toward understanding and compressing hybrid vision-language models for long videos. We introduce TimeViper, a Mamba-Transformer hybrid model equipped with TransV, an internal LLM token transfer module that compresses vision information into text tokens. We reveal that visual information gradually shifts into text tokens as depth of hybrid LLM layer increases, resulting in substantial redundancy among vision tokens in deep layers. TimeViper can efficiently process hour-long videos while maintaining strong multimodal understanding. TimeViper achieves promising performance for long video understanding across multi-choice video QA, temporal video grounding, and detailed video captioning.
Appendix
Appendix A Limitations
@@
First, our current performance still falls short of the SOTA models due to limited training data and insufficient model training. Second, while TransV enables processing over 10,000 frames, the model has not been trained on videos of such duration.
Appendix B Experimental Setups
Implementation details. When training attention-based dropping in multi-turn dialogue scenarios, the attention distribution is computed using the last token of the final instruction as the query. For temporal video grounding data, we incorporate a time-aware prompt [li2024videochatflash]: “The video lasts for {} seconds, and {} frames are uniformly sampled from it.” During training, we randomly sample one instruction from a pool of 15 manually constructed task prompts, such as: “From the video, locate the portion that aligns with the textual query , and output the start and end timestamps in seconds. The output format of the predicted timestamp should be like: ’start to end’ seconds. A specific example is : 12.0 to 20.0 seconds”. We do not use a system prompt, but we retain the BOS token to act as an attention sink [xiao2024attentionsink].
Training data summary. Our training pipeline adopts a two-stage strategy. We summarize the training data in Table 3. Specifically, in the first image-text alignment stage, we utilize 3 million images randomly sampled from the CC12M dataset [changpinyo2021cc12m], paired with captions sourced from PixelProse [singla2024pixelprose]. In the second video instruction tuning stage, we assemble a composite dataset to enhance MLLM’s video understanding and timestamp prediction capabilities, comprising: (1) 1.3M samples from LLaVA-Video [zhang2025llavavideo]; (2) 253K data from Kinetics400 [carreira2017k400] and WebVid [bain2021webvid] that are recaptioned with GPT-4o or Gemini by ShareGemini [sharegemini] and ShareGPT-4 [chen2024sharegpt4video]; (3) 100K samples from ET-Instruct [liu2024etbench]; (4) 112K samples from VideoGPT-Plus [Maaz2024VideoGPT+]; (5) 11K samples from LongVid [li2024videochatflash] and MovieChat [song2024moviechat]; (6) 26K dense video captioning (DVC) samples aggregated from ActivityNet [caba2015activitynet], COIN [tang2019coin], HiREST [zala2023hirest], ViTT [huang2020vitt], and YouCook2 [zhou2018youcook2]; and (7) 250K temporal video grounding (TVG) samples [wang2025timer1] from YT-Temporal [yang2023vid2seq], DiDeMo [anne2017didemo], QuerYD [oncescu2021queryd], InternVid [wang2024internvid], and HowTo100M [miech2019howto100m]. We obtain grounding data with annotations from VTG-IT [guo2025vtgllm], TimeIT [ren2024timechat], TimePro [zeng2025timesuite], HTStep [afouras2023htstep], and LongVid [li2024videochatflash]. This data collection process yields 339K temporal grounding samples. To ensure data quality, we apply a simple cleaning protocol to the TVG data. Specifically, we filter out coarse-grained samples where the ground truth duration exceeds 30 seconds or spans more than one-third of the total video length. We also discard invalid entries containing out-of-bound timestamps. Consequently, our TVG training data remains 250K.
Table 3: Data recipe. Overview of the datasets used in our two-stage training pipeline.
Stage 1: Projector Alignment
Image caption data (3M) CC12M (3M) [changpinyo2021cc12m] with PixelProse captions [singla2024pixelprose]
Stage 2: Video Instruction-Tuning
Image instruction data (2.8M) LLaVA-OneVision (2.8M) [lillavaov];
Video instruction data (1.8M) LLaVA-Video (1.3M) [zhang2025llavavideo]; Kinetics400 & WebVid (253K) [carreira2017k400, bain2021webvid] (recaptioned via ShareGemini [sharegemini] & ShareGPT-4 [chen2024sharegpt4video]); VideoGPT-Plus (112K) [Maaz2024VideoGPT+]; ET-Instruct (100K) [liu2024etbench]; LongVid [li2024videochatflash] & MovieChat (11K) [song2024moviechat]
Dense video captioning (26K) ActivityNet [caba2015activitynet], COIN [tang2019coin], HiREST [zala2023hirest], ViTT [huang2020vitt], YouCook2 [zhou2018youcook2]
Temporal video grounding (250K) YT-Temporal [yang2023vid2seq], DiDeMo [anne2017didemo], QuerYD [oncescu2021queryd], InternVid [wang2024internvid], HowTo100M [miech2019howto100m] (Annotated by VTG-IT [guo2025vtgllm], TimeIT [ren2024timechat], TimePro [zeng2025timesuite], HTStep [afouras2023htstep], LongVid [li2024videochatflash])
Table 4: Performance of applying TransV to Qwen2.5 and Nano.
avg.acc val M-Avg overall long avg.acc mIoU avg.acc
Qwen2.5-7B (ours w/ TransV) Qwen2.5-7B [qwen2.5] 55.7 53.7 63.3 55.7 47.4 37.0 38.7 40.7
TimeViper (ours w/ TransV) Nanov2-9B [basant2025nanov2] 56.2 52.0 63.1 56.9 48.2 35.6 37.9 39.1
Appendix C Main Results
TransV can be also applied to Qwen2.5. Qwen can also use TransV to handle ultra-long sequences, as shown in Table 4. We observe that on LVBench, this even brings a +0.4 improvement. However, the performance drop on VDC is more severe for Qwen than for Nano. For example, Qwen drops from 42.0 to 40.7 (a decrease of 1.3 points), whereas Nano drops from 39.7 to 39.1 (a decrease of 0.6 points).
Appendix D Qualitative Results
We define average attention scores used in both Mamba-2 and self-attention layers to analyze attentions received by different types of tokens.
Average attention score computation. We adopt the category-level attention score definition from LLaVA-Mini [zhang2025llavamini]. Tokens are grouped into instruction, vision, and response categories: $\mathcal{T}_{\text{ins}}$, $\mathcal{T}_{\text{vis}}$, and $\mathcal{T}_{\text{res}}$. Let $a_{ij}$ denote the attention score from token $t_{i}$ to token $t_{j}$, averaged over all attention heads. For two token categories $\mathcal{A},\mathcal{B}\in\{\mathcal{T}_{\text{ins}},\mathcal{T}_{\text{vis}},\mathcal{T}_{\text{res}}\}$, we define their category-level attention score as:
$\text{Attn}(\mathcal{A}\!\to\!\mathcal{B})=\frac{\sum_{t_{i}\in\mathcal{A}}\sum_{t_{j}\in\mathcal{B}}a_{ij}}{\left|\left\{\,t_{i}\in\mathcal{A}\;\big|\;\sum_{t_{j}\in\mathcal{B}}a_{ij}>0\right\}\right|}.$ (9)
The denominator counts the number of tokens in $\mathcal{A}$ that attend to any token in $\mathcal{B}$ with non-zero weight, ensuring that tokens masked by the causal attention mask are excluded.
In Figure 10, we analyze the overall attention scores from the entire sequence $\mathcal{T}=\mathcal{T}_{\text{ins}}\cup\mathcal{T}_{\text{vis}}\cup\mathcal{T}_{\text{res}}$ to a target category $\mathcal{B}$. To ensure equal contribution from each category, we compute the arithmetic mean of the scores:
$\text{Attn}(\mathcal{T}\!\to\!\mathcal{B})=\frac{1}{3}\big($ $\text{Attn}(\mathcal{T}_{\text{ins}}\!\to\!\mathcal{B})$ (10)
$+$ $\text{Attn}(\mathcal{T}_{\text{vis}}\!\to\!\mathcal{B})$
$+$ $\text{Attn}(\mathcal{T}_{\text{res}}\!\to\!\mathcal{B})\big).$
Qualitative results on VideoMME.  Figure 11(c) provides qualitative examples illustrating the effectiveness of TransV on the MCQ task. In the first case (top row), the query requires retrieving fine-grained visual information, specifically, the defending layers about the Berlin Wall. The compressed model, TimeViper w/ TransV, successfully attends to the critical frame at 03:36, which clearly depicts the structure of the wall, enabling it to select the correct answer. This example highlights TransV’s capability to identify and attend to key visual cues within long videos. In the second case (bottom row), the query requires long-term temporal reasoning to infer the chronological order of topics in a biology lecture. The correct answer relies on aggregating information across the entire video duration. In this case, a model must accurately align the textual concepts (e.g., structure, photosynthesis) with their corresponding temporal segments (00:52, 07:37, etc.), As shown in the visual evidence, TimeViper w/ TransV correctly deduces the sequence (c)-(d)-(e)-(a)-(b), demonstrating its capability in modeling global temporal dependencies and understanding narrative structure.
Qualitative results on Charades. For temporal video grounding results shown in Figure 11(b), we observe that incorporating the compression module yields only minimal changes. Both the original and compressed models correctly interpret timestamps and successfully localize video segments that correspond to the natural language query.
Qualitative results on VDC.  Figure 11(a) presents the qualitative comparison for the video detailed captioning task. In the generated captions, green text denotes accurate visual details, while red text indicates hallucinations or factual errors. In the first example showing a person painting (top row), the baseline model suffers from severe object hallucination, fabricating elements such as a “sponge” which are absent in the video. Surprisingly, TimeViper w/ TransV generates more faithful description, accurately recognizing specific objects like “paintbrushes”. This suggests that compression may help reduce hallucination by filtering out irrelevant or misleading visual information. In the second example (bottom row), the compression module TransV largely retains the original model behavior where there are both correct and incorrect captions.
(a) Qualitative results on VDC.
(b) Qualitative results on Charades.
(c) Qualitative results on VideoMME.
Figure 11: Qualitative results on three benchmarks.